OpenStudy (anonymous):

find the function of the form y=sin bx given the period. Pi over 2

OpenStudy (anonymous):

The normal period of \(\sin(x)\) is \(2\pi\).

OpenStudy (anonymous):

Now if you need to get to \(x=2\pi\) for the cycle to repeat, then were you to have \(\sin(bx)\) then you need to get to \(bx=2\pi\) for the cycle to repeat. Thus the period is \(2\pi/b\).

OpenStudy (anonymous):

So our equation here is \[ \frac{2\pi }{b}=\frac \pi 2 \]

OpenStudy (anonymous):

Solve for \(b\) and you're pretty much done.

OpenStudy (anonymous):

thanks wio